The intrinsic temperature of a semiconductor is the temperatures at which the intrinsic


carrier concentration equals the impurity concentration. Find the intrinsic temperature for a


silicon sample doped with 1015 phosphorus atoms/cm3


.

1
Expert's answer
2022-02-09T16:15:15-0500

"n=(\frac{M_1}{M_2})^{\frac 32}(\frac T{300})^\frac 32\cdot 2.5\cdot 10^{19},\implies"

"T=41~K ."